Circular No. 5762

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 R. G. Strom, Netherlands Foundation for Research in Astronomy

(NFRA), Dwingeloo, reports on behalf of the NFRA supernova team (A. J. Boonstra, R. Braun, A. G. de Bruyn, A. R. Foley, and himself): “We have made recent 6-cm observations of SN 1993J with the Westerbork Telescope. On Apr. 8.60-9.10 UT, no object brighter than 0.4 ± 0.2 mJy is seen within 10′′ of SN 1993J. In an observation on Apr. 12.59-13.09, we have possibly detected a source within 1′′ of the VLA position (IAUC 5759). Its flux density at 6 cm (4869 MHz) was 0.44 ± 0.16 mJy. This is consistent with detections made at shorter wavelengths (IAUC 5751, 5752, 5759), and also with the temporal development of the source spectrum. The flux density of the nuclear component in NGC 3031 is 123 mJy in both measurements.”

N. Panagia, Space Telescope Science Institute and University of Catania; and S. D. Van Dyk and K. W. Weiler, Naval Research Laboratory, write: “The available radio observations of SN 1993J are already sufficient for initial modelling of the progenitor star and projection of the likely evolution of the radio emission. Applying the Chevalier mini-shell model, which has been shown appropriate for other radio supernovae (Weiler et al. 1986, Ap.J. 301, 790) with the usual assumptions for the stellar wind from the red supergiant progenitor star, and assigning a normal spectral index of –0.7 and an index for the emission decay rate of –0.7, the model yields a pre-supernova mass loss rate of ∼ 2 × 10^{−6} M_{⊙}/yr. The stellar evolution models of Maeder and Meynet (1988, A.Ap. Suppl. 76, 411) then imply a ZAMS mass for the progenitor of SN 1993J of < 15 M_{⊙}. Comparison with other known radio supernovae, however, indicates that SN 1993J is quite similar in radio properties to SN 1980K and SN 1981K and may represent the lower end of the mass range of stars that can explode as type-II supernovae. This relative consistency of properties may indicate a relatively firm lower limit to the radio properties of type-II supernovae, and therefore a possible minimum distance indicator. Our modelling also results in predictions of the radio flux density to be expected from SN 1993J at various frequencies to assist planning observations at other telescopes. Taking Mar. 26 as day 1, the expected peak flux densities at different radio frequencies are as follows: Apr. 14 (day 20), ∼ 20 mJy at 23 GHz; Apr. 19, ∼ 20 mJy at 15 GHz; May 4, ∼ 25 mJy at 8.4 GHz; May 19, ∼ 30 mJy at 5 GHz; Aug. 2 (day 130), ∼ 35 mJy at 1.4 GHz. After peak, the flux density at each frequency should decline approximately as t^{−0.7}.”

1993 April 16 Daniel W. E. Green